OpenStudy (anonymous):

Choose the equivalent system of linear equations that will produce the same solution as the one given below.

OpenStudy (anonymous):

Choose the equivalent system of linear equations that will produce the same solution as the one given below. 4x - y = -11 2x + 3y = 5 (1 point) Question 5 options: 1) -4x - 9y = -19 -10y = -30 2) 4x + 3y = 5 2y = -6 3) 7x - 3y = -11 9x = -6 4) 12x - 3y = -33 14x = -28 Save

OpenStudy (imstuck):

The solutions for your solution are x = -2 and y = 3. We need to find out now which of your other systems give that same set of coordinates. Easy way to do that is to just plug in a -2 for x in one and see if it makes the equation true. If it doesn't move on to the next set. If it does, then you have to try the y value and make sure that one is true, also.

OpenStudy (imstuck):

-4(-2)-9(3)=-19? 8-27=-19? yes it does. So it works here. Let's try the other equation with this one. -10(3)=-30? yes it does. So this is the other system with the solutions of x = -2 and y = 3.

OpenStudy (imstuck):

4(-2)+3(3)=5? -8+9=5? no so it's not that one (i'm just going through to make sure there are not 2 sets of systems with the same solution set)

OpenStudy (imstuck):

7(-2)-3(3)=-11? -14-9=-11? no

OpenStudy (imstuck):

12(-2)-3(3)=-33? -24-9=-33 yes so let's try the other equation with this one. 14(-2)=-28 -28=-28 yes.

OpenStudy (imstuck):

So your answers are A and D.